# A probabilistic, data-driven closure model for RANS simulations with aleatoric, model uncertainty

### v1

@@ 2.2.3 Data, likelihood and posterior
The probabilistic model proposed is trained with indirect observational data that pertain to time-averaged velocities and pressures at various points in the problem domain. This is in contrast to the majority of efforts in data-driven RANS closure modeling (Geneva and Zabaras (2019); Ling et al. (2016b); Wang et al. (2017b); Kaandorp and Dwight (2020)), which employ direct RS data. In the ensuing numerical illustrations, the data is obtained from higher-fidelity computational simulations, but one could readily make use of actual, experimental observations.
@@
$\begin{array}[]{ll}p(\mathcal{D}\mid\bm{\theta},\bm{\epsilon}_{\tau}^{(1:M)})&=\prod_{m=1}^{M}p(\hat{\bm{z}}^{(m)}\mid\bm{\theta},\bm{\epsilon}_{\tau}^{(m)})\\ &=\prod_{m=1}^{M}\mathcal{N}(\hat{\bm{z}}^{(m)}\mid\bm{z}(\bm{\theta},\bm{\epsilon}_{\tau}^{(m)}),\bm{\Sigma})\end{array}$ (26)
where $\bm{z}(\bm{\theta},\bm{\epsilon}_{\tau}^{(m)})$ denotes the solution vector of the discretized RANS equations (see Equation (7)) with the closure model suggested by Equation (13) i.e. $\bm{\tau}=\bm{\tau}_{\bm{\theta}}(\bm{u})+\bm{\epsilon}_{\tau}^{(m)}$. We note that a different set of latent variables $\bm{\epsilon}_{\tau}^{(m)}$ is needed for each flow scenario as, by its nature, model discrepancy will in general assume different values under different flow conditions. We denote with $\bm{\Sigma}$ the covariance matrix which, given the absence of actual observation noise, plays the role of a tolerance parameter determining the tightness of the fit. The covariance was expressed as $\bm{\Sigma}=diag(\sigma_{1}^{2},\cdots,\sigma_{3N_{obs}}^{2})$, where the values in the diagonal vector are set to $1\%$ of the mean of the squares of each observable across the $M$ flow settings i.e. $\sigma_{i}^{2}=0.01~{}\frac{1}{M}\sum_{m=1}^{M}\left(\hat{z}_{i}^{(m)}\right)^{2}$.
@@
  • the hyperparameters $\bm{\Lambda}$ in the hyperprior of $\bm{E}_{\tau}^{(1:M)}$
@@
$\begin{array}[]{ll}p(\bm{\theta},\bm{E}_{\tau}^{(1:M)},\bm{\Lambda}|~{}\mathcal{D})&\propto p(\mathcal{D}~{}|~{}\bm{\theta},\bm{E}_{\tau}^{(1:M)})~{}p(\bm{E}_{\tau}^{(1:M)}|\bm{\Lambda})~{}p(\bm{\theta})~{}p(\bm{\Lambda})\\ &=\left(\prod_{m=1}^{M}p(\hat{\bm{z}}^{(m)}\mid\bm{\theta},\bm{E}_{\tau}^{(m)})~{}p(\bm{E}_{\tau}^{(m)}|\bm{\Lambda})\right)~{}p(\bm{\theta})~{}p(\bm{\Lambda})\end{array}$ (27)
@@ 2.2.4 Inference and Learning
On the basis of the probabilistic model proposed and the posterior formulated in the previous section, we discuss numerical inference strategies for identifying the unknown parameters and latent variables. The intractability of the posterior stems from the likelihood which entails the solution of the discretized RANS equations. We advocate the use of Stochastic Variational Inference (SVI) Hoffman et al. (2013) which results in a closed-form approximation of the posterior $p(\bm{\theta},\bm{E}_{\tau}^{(1:M)},\bm{\Lambda}|\mathcal{D})$. In contrast to the popular, sampling-based strategies (MCMC, SMC etc.), SVI yields biased estimates at the benefit of computational efficiency. Given a family of probability densities $q_{\bm{\xi}}\left(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)}\right)$ parametrized by $\bm{\xi}$, we find the optimal, i.e. the one that is closest to the exact posterior in terms of their Kullback-Leibler divergence, by maximizing the Evidence Lower Bound (ELBO) $\mathcal{F}(\bm{\xi})$ Bishop and Nasrabadi (2006):
@@
$=\mathbb{E}_{q_{\xi}\left(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)}\right)}\left[\log\left(\frac{p(\mathcal{D}~{}|~{}\bm{\theta},\bm{E}_{\tau}^{(1:M)})~{}p(\bm{E}_{\tau}^{(1:M)}|\bm{\Lambda})~{}p(\bm{\theta})~{}p(\bm{\Lambda})}{q_{\xi}\left(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)}\right)}\right)\right]$ (28)
As its name suggests, it can be readily shown that ELBO lower-bounds the model log-evidence and their difference is given by the aforementioned KL-divergence i.e.:
$\log p(\mathcal{D})=\mathcal{F}(\bm{\xi})+KL\left(q_{\xi}(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)})~{}\middle|\middle|~{}p(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)}|\mathcal{D})\right)$ (29)
In order to expedite computations we employ a mean-field assumption Blei et al. (2017) according to which the approximate posterior is factorized as:
@@
$q_{\bm{\xi}}(\bm{\theta})=\delta(\bm{\theta}-\bm{\theta}_{MAP})$ (31)
$q_{\bm{\xi}}(\bm{\Lambda})=\delta(\bm{\Lambda}-\bm{\Lambda}_{MAP})$ (32)
In essence, we obtain point estimates for $\bm{\theta},\bm{\Lambda}$ which coincide with the Maximum-A-Posteriori (MAP) estimates. For the model discrepancy variables $\bm{E}_{\tau}^{(m)}$, we employ diagonal Gaussians given by:
$q_{\bm{\xi}}(\bm{E}_{\tau}^{(m)})=\mathcal{N}(\bm{E}_{\tau}^{(m)}\mid\bm{\mu}^{(m)}_{E},\text{diag}(\bm{\sigma}^{2,(m)}_{E})),\quad\forall i\in\{1,\ldots,M\}$ (33)
@@
$\bm{\xi}=\{\bm{\theta}_{MAP},\bm{\Lambda}_{MAP},\{\bm{\mu}^{(m)}_{E},\bm{\sigma}^{2,(m)}_{E}\}_{m=1}^{M}\}$ (34)
The updates of the parameters $\bm{\xi}$ are carried out using derivatives of the ELBO. These entail expectations with respect to $q_{\bm{\xi}}$ which are estimated (with noise) by Monte Carlo in conjunction with the ADAM stochastic optimization scheme Kingma and Ba (2014). In order to reduce the Monte-Carlo noise in the estimates, we employ the reparametrization trick Kingma and Welling (2013). This is made possible here for two reasons: Firstly due to the form of the approximate posterior $q_{\bm{\xi}}$. In particular, if we summarily denote with $\bm{\eta}=\{\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)}\}$ and given that the approximate posterior $q_{\bm{\xi}}(\bm{\eta})$ can be represented by deterministic transform $\bm{\eta}=g_{\xi}(\bm{\phi})$, where $\bm{\phi}$ follows a known density $q(\bm{\phi})$222 Based on the form of $q_{\bm{\xi}}$ in Equations (30),(31), (32), (33), the transform employed can be written as $\bm{\theta}=\bm{\theta}_{MAP}$, $\bm{\Lambda}=\bm{\Lambda}_{MAP}$ and $\bm{E}_{\tau}^{(m)}=\bm{\mu}^{(m)}_{E}+\text{diag}(\bm{\sigma}^{2,(m)}_{E})~{}\bm{\phi}$ where $q(\bm{\phi})=\mathcal{N}(\bm{\phi}|~{}\bm{0},\bm{I})$. , the expectations involved in the ELBO and, more importantly, in its gradient can be rewritten as:
$\nabla_{\bm{\xi}}\mathcal{F}(\bm{\xi})$ $=\mathbb{E}_{q(\bm{\phi})}\left[\nabla_{\bm{\xi}}g_{\xi}(\bm{\phi})\nabla_{\bm{\eta}}\left(\log p(\mathcal{D},\bm{\eta})-\log q_{\xi}(\bm{\eta})\right)\right]$ (35)
@@
Further details about the derivatives of the ELBO and the use of RANS-model sensitivities obtained by an adjoint formulation can be found in B.
An algorithmic summary of the steps entailed is contained in Algorithm 1.
@@
while ELBO $-^\hat{\mathcal{F}}$ not converged do
@@
       Estimate gradient of ELBO $\hat{\nabla_{\bm{\xi}}\mathcal{F}}$ using Equation (53) ;
@@
       $\bm{\xi}^{(n+1)}\leftarrow\bm{\xi}^{(n)}+\bm{\rho}^{(n)}\odot\hat{\nabla_{\bm{\xi}}\mathcal{F}}$ ;
@@ 2.2.5 Predictions
In this section, we describe how probabilistic, predictive estimates of any quantity of interest related to the RANS-simulated flow can be produced using the trained model. In particular, one can obtain a predictive, posterior density $p(\mathbf{z}|\mathcal{D})$ on the whole solution vector $\mathbf{z}$ of the RANS equations as follows:
@@
The third of the densities in the intergrand is the posterior which is substituted by its variational approximation i.e. $q_{\bm{\xi}}$ in Equation (30) and for the optimal parameter values $\bm{\xi}$ identified as described in the previous section. The second of the densities represents the prior model prescribed in Equation (24). Finally the first of the densities is simply a Dirac-delta that corresponds to the solution of the RANS equations obtained when using a closure model of the form $\bm{\tau}=\tau_{\bm{\theta}}(\bm{u})+\bm{W}\bm{E}_{\tau}$. In practical terms and given the intractability of this integral, the equation above suggests a Monte Carlo scheme for obtaining samples from $p(\bm{z}|\bm{D})$ which involves the following steps. For each sample:
@@
The aforementioned steps would need to be repeated for as many samples as desired. Subsequently the samples can be used to compute statistics of the predictive estimate (e.g. predictive mean, variance, credible intervals etc) not only of $\mathbf{z}$ (i.e. velocities/pressures) but of any quantity of interest such as the lift, drag, skin friction etc. We note however that stochastic RS discrepancy terms $\bm{\epsilon}_{\tau}$ or $\bm{E}_{\tau}$ and the associated probabilistic model, are limited to the flow geometry used for the training. While it can be used for unseen flow scenarios (e.g. different $Re$ number, inlet conditions, boundary conditions), it cannot be employed to a different flow geometry. In theory, the parametrization of the $\bm{\epsilon}_{\tau}$ can be updated to accommodate different geometries, but we leave it for future investigations. Finally, we would like to highlight the fact that baseline RANS data is not needed as an input to the neural networks for prediction in the proposed scheme, as opposed to other frameworks that have been employed in the past (e.g. Geneva and Zabaras (2019); Ling et al. (2016b); Kaandorp and Dwight (2020); Wang et al. (2017b)).
In terms of computational aspects, the stochastic nature of the (reduced) discrepancy tensor $\bm{E}_{\tau}$ can potentially introduce non-smoothness in the RS vector $\bm{\tau}$ used for solving the RANS equations. In the present work, we have used diagonal covariance for the hyper-prior of $\bm{E}_{\tau}$ given by $\bm{\Lambda}=diag(\bm{\Lambda}^{(J)}),~{}~{}J=1,\ldots,N_{d}$, thus assuming there is no correlation among the nearby nodes/region. As an avenue for future work, a banded covariance matrix can be employed to capture such spatial correlations. Sparsity-inducing priors that account for spatial correlations have been proposed in Bardsley (2013); Wu et al. (2014). Numerical results, training data and the corresponding source code will be made available at https://github.com/pkmtum/D3C-UQ/ upon publication.

### v2

@@ 2.2.3 Data, likelihood and posterior
The probabilistic model proposed is trained with indirect observational data that pertain to time-averaged velocities and pressures at various points in the problem domain. This is in contrast to the majority of efforts in data-driven RANS closure modeling [22, 15, 65, 17], which employ direct RS data. In the ensuing numerical illustrations, the data is obtained from higher-fidelity computational simulations, but one could readily make use of actual, experimental observations.
@@
$\begin{array}[]{ll}p(\mathcal{D}\mid\bm{\theta},\bm{\epsilon}_{\tau}^{(1:M)})&=\prod_{m=1}^{M}p(\hat{\bm{z}}^{(m)}\mid\bm{\theta},\bm{\epsilon}_{\tau}^{(m)})\\ &=\prod_{m=1}^{M}\mathcal{N}(\hat{\bm{z}}^{(m)}\mid\bm{z}(\bm{\theta},\bm{\epsilon}_{\tau}^{(m)}),\bm{\Sigma}),\end{array}$ (22)
where $\bm{z}(\bm{\theta},\bm{\epsilon}_{\tau}^{(m)})$ denotes the solution vector of the discretized RANS equations (see Equation (7)) with the closure model suggested by Equation (13) i.e. $\bm{\tau}=\bm{\tau}_{\bm{\theta}}(\bm{u})+\bm{\epsilon}_{\tau}^{(m)}$. We note that a different set of latent variables $\bm{\epsilon}_{\tau}^{(m)}$ is needed for each flow scenario as, by its nature, model discrepancy will in general assume different values under different flow conditions. We also note that the $\bm{\epsilon}_{\tau}$ is grid-dependent or problem geometry-dependent, which restricts it from making predictions for a completely different flow geometry. One could, with appropriate modeling enhancements, learn the dependence of $\bm{\epsilon}_{\tau}$ on flow parameters (e.g. boundary/initial conditions, geometry, $Re$ number) which would make it usable in other flow geometries. Alternatively, it could be trained under different flow settings (e.g. different $Re$ numbers as we do) and learn, on aggregate, the model error which in turn could be used to make predictions under different flow settings (as in our case for different $Re$ number).
We denote with $\bm{\Sigma}$ the covariance matrix of the Gaussian likelihood which, given the absence of actual observation noise, plays the role of a tolerance parameter determining the tightness of the fit. The covariance was expressed as $\bm{\Sigma}=diag(\sigma_{1}^{2},\cdots,\sigma_{3N_{obs}}^{2})$, where the values in the diagonal vector are set to $1\%$ of the mean of the squares of each observable across the $M$ flow settings i.e. $\sigma_{i}^{2}=0.01~{}\frac{1}{M}\sum_{m=1}^{M}\left(\hat{z}_{i}^{(m)}\right)^{2}$.
@@
  • the hyperparameters $\bm{\Lambda}$ in the hyperprior of $\bm{E}_{\tau}^{(1:M)}$,
@@
$\begin{array}[]{ll}p(\bm{\theta},\bm{E}_{\tau}^{(1:M)},\bm{\Lambda}|~{}\mathcal{D})&\propto p(\mathcal{D}~{}|~{}\bm{\theta},\bm{E}_{\tau}^{(1:M)})~{}p(\bm{E}_{\tau}^{(1:M)}|\bm{\Lambda})~{}p(\bm{\theta})~{}p(\bm{\Lambda})\\ &=\left(\prod_{m=1}^{M}p(\hat{\bm{z}}^{(m)}\mid\bm{\theta},\bm{E}_{\tau}^{(m)})~{}p(\bm{E}_{\tau}^{(m)}|\bm{\Lambda})\right)~{}p(\bm{\theta})~{}p(\bm{\Lambda}).\end{array}$ (23)
@@ 2.2.4 Inference and Learning
On the basis of the probabilistic model proposed and the posterior formulated in the previous section, we discuss numerical inference strategies for identifying the unknown parameters and latent variables. The intractability of the posterior stems from the likelihood which entails the solution of the discretized RANS equations. We advocate the use of Stochastic Variational Inference (SVI) [76] which results in a closed-form approximation of the posterior $p(\bm{\theta},\bm{E}_{\tau}^{(1:M)},\bm{\Lambda}|\mathcal{D})$. In contrast to the popular, sampling-based strategies (MCMC, SMC etc.), SVI yields biased estimates at the benefit of computational efficiency. Readers interested in comparative studies/discussions on SVI and MCMC in terms of accuracy and computational efficiency are directed to [77]. Given a family of probability densities $q_{\bm{\xi}}\left(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)}\right)$ parametrized by $\bm{\xi}$, we find the optimal, i.e. the one that is closest to the exact posterior in terms of their Kullback-Leibler divergence, by maximizing the Evidence Lower Bound (ELBO) $\mathcal{F}(\bm{\xi})$ [68]:
@@
$=\mathbb{E}_{q_{\xi}\left(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)}\right)}\left[\log\left(\frac{p(\mathcal{D}~{}|~{}\bm{\theta},\bm{E}_{\tau}^{(1:M)})~{}p(\bm{E}_{\tau}^{(1:M)}|\bm{\Lambda})~{}p(\bm{\theta})~{}p(\bm{\Lambda})}{q_{\xi}\left(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)}\right)}\right)\right].$ (24)
As its name suggests, it can be readily shown and that ELBO lower-bounds the model log-evidence and their difference is given by the aforementioned KL-divergence i.e.:
$\log p(\mathcal{D})=\mathcal{F}(\bm{\xi})+KL\left(q_{\xi}(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)})~{}\middle|\middle|~{}p(\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)}|\mathcal{D})\right).$ (25)
In order to expedite computations we employ a mean-field assumption [77] according to which the approximate posterior is factorized as:
@@
$q_{\bm{\xi}}(\bm{\theta})=\delta(\bm{\theta}-\bm{\theta}_{MAP}),$ (27)
$q_{\bm{\xi}}(\bm{\Lambda})=\delta(\bm{\Lambda}-\bm{\Lambda}_{MAP}).$ (28)
In essence, we obtain point estimates for $\bm{\theta},\bm{\Lambda}$ which coincide with the Maximum-A-Posteriori (MAP) estimates. The specific choice of Dirac-deltas was motivated by computational cost reasons in the inference and prediction steps. Furthermore, given that the dimension of $\bm{\theta}$ is generally (much) smaller than the number of observations (see section 3), we anticipate that its posterior would not deviate much from a Dirac-delta. The final reason for which posterior uncertainty (albeit small) on $\bm{\theta}$ was not included, was to emphasize the significance and impact of the stochastic, model correction term that we propose.
For the model discrepancy variables $\bm{E}_{\tau}^{(m)}$, we employ uncorrelated Gaussians given by:
$q_{\bm{\xi}}(\bm{E}_{\tau}^{(m)})=\mathcal{N}(\bm{E}_{\tau}^{(m)}\mid\bm{\mu}^{(m)}_{E},\text{diag}(\bm{\sigma}^{2,(m)}_{E})),\quad\forall i\in\{1,\ldots,M\}.$ (29)
@@
$\bm{\xi}=\{\bm{\theta}_{MAP},\bm{\Lambda}_{MAP},\{\bm{\mu}^{(m)}_{E},\bm{\sigma}^{2,(m)}_{E}\}_{m=1}^{M}\}.$ (30)
The updates of the parameters $\bm{\xi}$ are carried out using derivatives of the ELBO. These entail expectations with respect to $q_{\bm{\xi}}$ which are estimated (with noise) by Monte Carlo in conjunction with the ADAM stochastic optimization scheme [78]. In order to reduce the Monte-Carlo noise in the estimates, we employ the reparametrization trick [79]. This is made possible here due to the form of the approximate posterior $q_{\bm{\xi}}$. In particular, if we summarily denote with $\bm{\eta}=\{\bm{\theta},\bm{\Lambda},\bm{E}_{\tau}^{(1:M)}\}$ and given that the approximate posterior $q_{\bm{\xi}}(\bm{\eta})$ can be represented by deterministic transform $\bm{\eta}=g_{\xi}(\bm{\phi})$, where $\bm{\phi}$ follows a known density $q(\bm{\phi})$222 Based on the form of $q_{\bm{\xi}}$ in Equations (26),(27), (28), (29), the transform employed can be written as $\bm{\theta}=\bm{\theta}_{MAP}$, $\bm{\Lambda}=\bm{\Lambda}_{MAP}$ and $\bm{E}_{\tau}^{(m)}=\bm{\mu}^{(m)}_{E}+\text{diag}(\bm{\sigma}^{2,(m)}_{E})~{}\bm{\phi}$ where $q(\bm{\phi})=\mathcal{N}(\bm{\phi}|~{}\bm{0},\bm{I})$. , the expectations involved in the ELBO and, more importantly, in its gradient can be rewritten as:
$\nabla_{\bm{\xi}}\mathcal{F}(\bm{\xi})$ $=\mathbb{E}_{q(\bm{\phi})}\left[\nabla_{\bm{\xi}}g_{\xi}(\bm{\phi})\nabla_{\bm{\eta}}\left(\log p(\mathcal{D},\bm{\eta})-\log q_{\xi}(\bm{\eta})\right)\right].$ (31)
@@
Further details about the derivatives of the ELBO and the use of RANS-model sensitivities obtained by an adjoint formulation can be found in B. An algorithmic summary of the steps entailed is contained in Algorithm 1. We note finally that the ELBO $\mathcal{F}$ (which serves as a lower bound to the model evidence) and can be used to compare (in a Bayesian sense) alternative models. Such model variations could arise by changing e.g. $\bm{W}$ i.e. the partition into subdomains which enables the dimensionality reduction of the stochastic model discrepancy vector. One could even employ an adaptive division into subdomains guided by $\mathcal{F}$ so as the data can dictate which regions require more/less refinement.
@@
while ELBO $\hat{\mathcal{F}}$ not converged do
@@
       Estimate gradient of ELBO $\nabla_{\bm{\xi}}\mathcal{F}$ using Equation (45) ;
@@
       $\bm{\xi}^{(n+1)}\leftarrow\bm{\xi}^{(n)}+\bm{\rho}^{(n)}\odot\nabla_{\bm{\xi}}\mathcal{F}$ ;
@@ 2.2.5 Predictions
In this section, we describe how probabilistic, predictive estimates of any quantity of interest related to the RANS-simulated flow can be produced using the trained model. In particular, one can obtain a predictive, posterior density $p(\bm{z}|\mathcal{D})$ on the whole solution vector $\bm{z}$ of the RANS equations as follows:
@@
The third of the densities in the integrand is the posterior which is substituted by its variational approximation i.e. $q_{\bm{\xi}}$ in Equation (26) and for the optimal parameter values $\bm{\xi}$ identified as described in the previous section. The second of the densities represents the prior model prescribed in Equation (20). Finally, the first of the densities is simply a Dirac-delta which corresponds to the solution of the RANS equations obtained when using the proposed closure model for given $\bm{E}_{\tau},\bm{\theta}$. Since the RANS solver is a black-box, it would not be possible to propagate the uncertainty in $\bm{E}_{\tau}$ (and potentially $\bm{\theta}$) otherwise. In practical terms and given the intractability of this integral, the equation above suggests a Monte Carlo scheme for obtaining samples from $p(\bm{z}|\bm{D})$ which involves the following steps. For each sample:
@@
Other numerical integration techniques (e.g. Quasi Monte Carlo or Importance Sampling) could be used in order to obtain estimates of the integral with fewer RANS solves. The aforementioned steps would need to be repeated for as many samples as desired. Subsequently, the samples can be used to compute statistics of the predictive estimate (e.g. predictive mean, variance, credible intervals, etc) not only of $\bm{z}$ (i.e. velocities/pressures) but of any quantity of interest such as the lift, drag, skin friction, etc.
We note however that stochastic RS discrepancy terms $\bm{\epsilon}_{\tau}$ or $\bm{E}_{\tau}$ and the associated probabilistic model, are limited to the flow geometry used for the training. While it can be used for unseen flow scenarios (e.g. different $Re$ number, inlet conditions, boundary conditions), it cannot be employed for a different flow geometry. In theory, the parametrization of the $\bm{\epsilon}_{\tau}$ can be updated to accommodate different geometries, but we leave it for future investigations. Finally, we would like to highlight the fact that baseline RANS data is not needed as an input to the neural networks for prediction in the proposed scheme, as opposed to other frameworks that have been employed in the past (e.g. [22, 15, 17, 65]).
In terms of computational aspects, the stochastic nature of the (reduced) discrepancy tensor $\bm{E}_{\tau}$ can potentially introduce non-smoothness in the RS vector $\bm{\tau}$ used for solving the RANS equations. In the present work, we have used diagonal covariance for the hyper-prior of $\bm{E}_{\tau}$ given by $\bm{\Lambda}=diag(\bm{\Lambda}^{(J)}),~{}~{}J=1,\ldots,N_{d}$, thus assuming there is no correlation among the nearby nodes/region. As an avenue for future work, a banded covariance matrix can be employed to capture such spatial correlations. Sparsity-inducing priors that account for spatial correlations have been proposed in [80, 73]. A schematic overview of the methodological framework discussed in Section (2.2.4) and Section (2.2.5) is presented in Fig. (2). Numerical results, training data and the corresponding source code will be made available at https://github.com/pkmtum/D3C-UQ/ upon publication.
Figure 2: Schematic illustration of the training/inference (left block) and probabilistic prediction (right block) framework proposed.